OpenStudy (anonymous):

if C'(500)=75 and R'(500)=100, should the quantity produced be increased or decreased from q=500 in order to increase profits? if C'(500)=75 and R'(500)=100, should the quantity produced be increased or decreased from q=500 in order to increase profits? @Mathematics

OpenStudy (dumbcow):

to max profits R'(q) - C'(q) = 0 call this quantity Q if q<Q, then R'(q) - C'(q) >0 if q>Q, then R'(q) - C'(q) <0 for q=500 100 - 75 = 25 since 25 >0, assume 500 < Q therefore increase quantity to increase profits